Suppose Jane’s demand for a good is given by Jane consumes 120 units when Q=a - 6P. The price is $20. Suppose the price of the good increases by 5 percent to $21. What is the consumer surplus?
Answer #1

Q = a - - 6P %3D at @ =120 & p=20, we have (20 = a - 6X20 120 = a- 120 > a = 240 So, Demand Curve ! @= 240-6P 240 е чо 40 * N